1.Fairfield purchased the plant on March 1, 20X2, for $46,790,000. Additional costs to get it up and running were $3,780,000. Fairfield assigned a thirty-year useful life and residual value of $4,000,000 and used double-declining balance to depreciate the plant. Record the acquisition of the plant and depreciation for three years, assuming that Fairfield does not use the half-year convention.

2.On Dec 31, 20X4, Fairfield’s auditors raise concerns that the plant’s market value might be below its book value due to the failure of the j phone. They believe this decline is permanent and decide to test for impairment. The accountants and auditors agree that the plant will generate net cash flows of approximately $2,000,000 each year for the next fifteen years. Perform a test of recoverability on the plant.



1
Expert's answer
2021-07-19T15:40:17-0400
  1. Debit investments in non-current assets credit settlements with contractors - 46 790 000

Debit investments in non-current assets credit settlements with contractors - 3 780 000

Debit Fixed assets Credit investments in non-current assets - 50 570 000

annual depreciation rate:

"\frac{100}{30}=3,33"

Multiply the book value at the beginning of the period by double the usual annual rate:

"50 570 000\\times2\\times3.33=3\u00a0371\u00a0333.33"

Let's subtract the annual depreciation costs from the value of the initial period to calculate the value of the final period:

1 year: 50 570 000-3 371 333.3=47 198 666,67

2 year: "47 198 666.67\\times2\\times3.33=3\u00a0143\u00a0431.20"

47 198 666.67-3 143 431.20=44 055 235.47

3 year: "44 055 235.47\\times2\\times3.33=2\u00a0934\u00a0078.68"

44 055 235.47-2 934 078.68=41 121 156.79

2.If the expected cash flows exceed the net carrying amount of the asset, the asset is considered "recoverable" and no impairment is indicated.

"2 000 000\times15=30 000 000"

and residual value of the plant is $4,000,000 

In this case, the expected cash flows exceed the net carrying amount of the asset, the asset is considered "recoverable" and no impairment is indicated .